# Aspen Open Jets: Unlocking LHC Data for Foundation Models in Particle Physics

### v1

@@ Abstract
Foundation models are deep learning models pre-trained on large amounts of data which are capable of generalizing to multiple datasets and/or downstream tasks. This work demonstrates how data collected by the CMS experiment at the Large Hadron Collider can be useful in pre-training foundation models for HEP. Specifically, we introduce the AspenOpenJets dataset, consisting of approximately 180M high $p_{T}$ jets derived from CMS 2016 Open Data. We show how pre-training the OmniJet-$\alpha$  foundation model on AspenOpenJets improves performance on generative tasks with significant domain shift: generating boosted top and QCD jets from the simulated JetClass dataset. In addition to demonstrating the power of pre-training of a jet-based foundation model on actual proton-proton collision data, we provide the ML-ready derived AspenOpenJets dataset for further public use.
@@ I Introduction
While particle physics has long used machine learning techniques and is leading the way in adopting modern deep learning methods, the development and application of powerful foundation models – which have transformed natural language processing and computer vision – is still in its early stages in the field. A foundation model is a model that has been pre-trained on a large amount of data for a specific task, and that can be used for different tasks downstream Bommasani et al. (2022). The pre-training thus serves as a foundation, upon which the downstream tasks can be built. Well-known foundation models include large language models (LLMs) such as GPT-3 Brown et al. (2020), BERT Devlin et al. (2019) and LLaMA Touvron et al. (2023), image models such as DALL-E 2 Ramesh et al. (2022) and Imagen Saharia et al. (2022), and mixed-modality models such as CLIP Radford et al. (2021) and ALIGN Jia et al. (2021).
@@
There are several advantages to the use of foundation models. One example is the potential to boost the performance achievable on small datasets. Having access to a model pre-trained on a larger, similar dataset provides a ‘head start’ for the downstream task, allowing it to focus on the intricacies of the smaller dataset rather than having to rediscover the general structure of the data. In addition, a foundation model can save both human and computational resources. While pre-training may be a resource-intensive task, the downstream models would require less training, less data, and less time spent on optimization compared to what would be required if these models were built from scratch. Another interesting aspect is that a trained foundation model could enable implicit sharing of data and resources between different parts of the research community, a sharing that would not otherwise be possible.
@@
Foundation models are poised to play a central role in particle physics, especially in the context of experiments at the Large Hadron Collider (LHC). They are able to process the huge amounts of complex and diverse data generated by the LHC and have the potential to integrate multiple modalities - such as images, high-level event features or particle four-vectors - allowing more comprehensive data analyses. Current work on foundation models for particle physics has mainly focused on different event generation and classification tasks Kishimoto et al. (2023); Qu et al. (2022a); Heinrich et al. (2024); Birk et al. (2024); Harris et al. (2024); Mikuni and Nachman (2024), investigating pre-training strategies, encoding schemes and different architectures, and quantifying the transfer learning and generalization capabilities of these models.
@@
The LHC experiments have made significant amounts of data publicly available CMS Collaboration (2022a, b, 2024a, 2024b); ATLAS Collaboration (2024). Previous open data studies of CMS jets Larkoski et al. (2017); Tripathee et al. (2017); Komiske et al. (2020) have explored the substructure and energy flow of the jets. Foundation models can leverage the LHC open data for training and fine-tuning, potentially integrating different data modalities and data from different experiments. Furthermore, training foundation models on real data will result in a more accurate representation of particle physics processes than training on simulations. Finally, foundation models trained on a large variety of actual LHC data will help to promote transparency and collaboration in particle physics research.
@@
In this paper we present the AspenOpenJets (AOJ) dataset and study the transfer learning of foundation models obtained with this dataset. The AOJ dataset has been constructed from the CMS 2016 open data ‘JetHT’ CMS Collaboration (2024a, b) and prepared in a form that allows easy use of the dataset for phenomenological studies, especially for training machine learning models. With $180\,\rm M$ jets, AOJ is by far the largest dataset of its kind. We demonstrate an application of this dataset in a foundation model context, by pre-training the OmniJet-$\alpha$  Birk et al. (2024) foundation model on AOJ, which is expected to mostly contain QCD jets, and then fine-tuning it to generate jets from the JetClass Qu et al. (2022b) dataset of simulated jets. We do two such fine-tunings: first for JetClass QCD jets, which occupy a different kinematic region compared to the jets in AOJ, and then for top jets, which differ significantly in several features. We show that pre-training on AOJ indeed benefits the downstream task, making it easier for the model to generate these types of jets with fewer training examples compared to if it had been trained from scratch.
@@
The utility of CMS open data in a pre-training setting has been previously explored in Kishimoto et al. (2023), using 1M single-lepton events at the object level which were collected by CMS in 2015. Here we enlarge the scope and complexity by several orders of magnitude to encompass 180M jet events at the full constituent level. Furthermore, the authors of Kishimoto et al. (2023) used event selections to shape the open data into a similar topology to the intended target simulated data. In our work, no such shaping is needed.
@@ II Data
Figure 1: Comparison of AspenOpenJets (AOJ) and QCD JetClass distributions for particle and jet level observables. The last panel shows the average number of jet constituents for each particle species, accompanied by the corresponding $1\sigma$ standard deviation bands.
@@ II.1 CMS open data
The CMS experiment Chatrchyan et al. (2008) is a large general-purpose detector at the LHC. CMS has recently released 16.4 fb-1 of data from their 2016 runs (G and H) CMS Collaboration (2024a, b) and a corresponding set of simulations. In contrast to previous data releases CMS Collaboration (2014); CMS Collaboration (2022a, b), this data has been provided in the simplified MINIAOD format Petrucciani et al. (2015), as well as the even further simplified NANOAOD format Peruzzi et al. (2020). In these simplified analysis formats, the central CMS reconstruction steps have been applied, discarding much of the low-level content of the event in favor of a simplified description of the high-level objects (jets, electrons, muons, etc.) commonly used in data analysis. These formats are therefore significantly easier to use for those outside of the collaboration.
@@
CMS uses a particle-flow algorithm Sirunyan et al. (2017), which combines the information from the different detector systems to reconstruct and identify each individual particle in an event. Particles are classified as either muons, electrons, photons, charged hadrons or neutral hadrons. Jets are clustered from the particle-flow candidates in an event using the anti-$k_{t}$ jet finding algorithm Cacciari et al. (2008, 2012). The typical large-radius jet size used in CMS for jet substructure searches and measurements is $R=0.8$, commonly referred to as AK8 jets. For AK8 jets the pileup-per-particle identification algorithm (PUPPI) Sirunyan et al. (2020b); Bertolini et al. (2014) is used to mitigate the effect of pileup. The PUPPI algorithm assigns a weight to each particle based on the probability that it originates from the primary vertex or from pileup. The momentum of AK8 jets is determined from the weighted vector sum of all particles inside the jet. Additional corrections are applied based on simulation and data-driven measurements to better calibrate the measured jet momentum to the true momentum of the particle from the hard process Khachatryan et al. (2017b).
The NANOAOD format does not natively store information of the jet constituents. However, an extended version (PFNANO), which does include this information, can be easily produced from MINIAOD using tools provided by CMS Collaboration (2024). We therefore, for all data and simulations samples used, start with the MINIAOD format, process it to PFNANO and then apply our selections to select the jets of interest for our dataset.
In order to construct the largest sample of jets possible, we consider all events from the ‘JetHT’ CMS Collaboration (2024a, b) Run-2 dataset released by CMS to this date. The ‘JetHT’ dataset includes all events that have been selected by one of the HLT triggers related to jet momenta or total event hadronic activity (HT). Because we do not specify a specific trigger, the sample consists of events from an inclusive ‘or’ of any of the available triggers of this dataset. We note that this choice works well for building a large dataset for the purposes of foundation model training, but may introduce subtle kinematic biases resulting from the combined thresholds of the different triggers, which would complicate its use in a physics-based analysis.
@@
Applying these selections to the CMS open data yields a dataset of 178 million jets, which we call the AspenOpenJets (AOJ). Due to the large cross section of QCD processes, the vast majority of these jets are expected to originate from light quarks and gluons. A small fraction of the jets are expected to originate from the decays of boosted heavy resonances. A rough estimate of the number of boosted fully merged jets produced by W, Z, top, and Higgs bosons in the $b\bar{b}$ decay mode are estimated with Monte Carlo (MC) simulations provided by CMS CMS Collaboration (2024c), with the remainder assumed to originate from QCD. The numbers are reported in Table 1. More details on the MC samples used are given in Appendix A.
@@
For each jet we store its transverse momentum ($p_{T}$), pseudorapidity ($\eta$), and azimuthal angular coordinate ($\phi$). We also store its mass, groomed with the softdrop algorithm Larkoski et al. (2014) as computed within the CMS reconstruction. Up to 150 constituents of the jet are stored. For each constituent, its 4-momentum is stored in the format $(p_{x},p_{y},p_{z},E)$. We additionally store its transverse impact parameter ($d_{0}$) and longitudinal impact parameter ($d_{z}$) with their uncertainties, the charge of the candidate, its particle-ID (PID) in the PDG format [^2] Note that neutral hadrons $({\rm h}^{0})$ are assigned the ${\rm PID}=130$ of the neutral kaon $K_{L}^{0}$ while positively/negatively charged hadrons $({\rm h}^{\pm})$ are assigned ${\rm PID}=\pm 211$ of the charged pion. Navas et al. (2024) and its weight from the PUPPI algorithm. We also include additional jet substructure quantities computed within the CMS reconstruction, including the number of constituents in the jet, $N$-subjettiness variables Thaler and Van Tilburg (2011), various jet-tagging observables from the CMS implementation of ParticleNet Qu and Gouskos (2020); par (2020) and a regression of the jet mass from ParticleNet par (2020).
@@
II.2 JetClass
The JetClass Qu et al. (2022b) dataset is based on simulations and was first introduced in Qu et al. (2022a). It contains both particle-level and jet-level information for $125\,\rm M$ simulated jets, distributed over 10 different jet types and divided into a training ($100\,\rm M$), validation ($5\,\rm M$) and test ($20\,\rm M$) set. The simulation is performed with MadGraph5_aMC@NLO Alwall et al. (2014) for the hard process, followed by Pythia Sjöstrand et al. (2015) for parton showering and hadronization, and Delphes de Favereau et al. (2014) with the CMS detector Chatrchyan et al. (2008) card for the detector response. Jets are required to have a transverse momentum of $500\,\mathrm{GeV}<p_{T}<1000\,\mathrm{GeV}$ and a pseudorapidity $\eta<2$, which is more restrictive than the criteria used for AOJ. We overlay the JetClass QCD particle- and jet-level distributions (blue histograms) in fig. 1 to facilitate easy comparison with the corresponding AOJ distributions.
@@ III.1 Method
The foundation model chosen for this work is OmniJet-$\alpha$  Birk et al. (2024), a GPT-style Radford et al. (2018) model that predicts the next token in a sequence. The features of the jet constituents [^3] The jet constituents are $p_{T}$ ordered and up to the first 128 constituents are considered when training the models. are first converted into integer tokens using a VQ-VAE van den Oord et al. (2018); Bao et al. (2022); Heinrich et al. (2024); Huh et al. (2023) model, representing jets as sequences of tokens. The backbone of OmniJet-$\alpha$ is trained to predict the next token in these sequences. After pre-training, the model can generate new sequences autoregressively, which are then decoded by the frozen VQ-VAE back into physical space. Additionally, the pre-trained backbone weights can be loaded and used for other tasks, such as jet classification. In this work, we focus on dataset transfer learning for jet generation and leave the exploration of task switching for future studies.
@@
The features used for training are the constituent kinematics ($p_{T}$, $\Delta\eta$, $\Delta\phi$), where $\Delta\eta=\eta-\eta_{\text{jet}}$ and $\Delta\phi=\phi-\phi_{\text{jet}}$ represent the relative coordinates with respect to the jet axis. Before tokenization, these features were preprocessed using standardization. The AOJ dataset was tokenized using a codebook size of 8192 tokens, employing VQ-VAE hyperparameters as outlined in Birk et al. (2024), with two modifications: the number of attention heads was reduced from 8 to 4 without sacrificing performance, and the AdamW optimizer was replaced with the Ranger optimizer Wright (2019). A larger codebook with $\sim\!32\,\rm K$ tokens was also tested but produced results comparable to the $8\,\rm K$ configuration. Therefore, we only present results for the $8\,\rm K$ tokenization. The backbone model was trained on the full $180\,\rm M$ AOJ dataset with hyperparameters as specified in Birk et al. (2024). The model was trained for 900K steps on 12 GPUs with a batch size of 256 per GPU, and the final state at the end of the training was selected as the backbone model.
@@
To evaluate the transfer learning capabilities of OmniJet-$\alpha$ , we devised a downstream task to test its ability to generate jets of a different type than those it was pre-trained on, and to adapt to datasets with different kinematic cuts. The fine-tuning process involved loading the pre-trained backbone weights and retraining it using jets from the JetClass Qu et al. (2022b) dataset. To systematically evaluate the performance of the model, we trained it on six datasets with $10^{2}$, $10^{3}$, $10^{4}$, $10^{5}$, $10^{6}$, $10^{7}$ jets to study the effect of dataset size $D$, and fine-tuned the model separately on QCD jets ($q/g$) and top-quark jets ($t\to bqq^{\prime}$) to evaluate its adaptability to different jet substructure types. For each configuration, the model with the lowest validation loss during fine-tuning was selected for jet generation. To provide a baseline for comparison, we used a separate VQ-VAE tokenizer and trained generative models with the same backbone architecture from scratch using randomly initialized weights. This VQ-VAE is the one from Ref. Birk et al. (2024) that was trained on all 10 classes of the JetClass dataset. The generative models were trained on the tokenized QCD and top-quark samples. This approach allowed us to quantify the performance improvement achieved by pre-training on the AOJ dataset.
@@
Both training strategies used a batch size of 256444For training samples of 100 jets, the batch size was reduced to 100. per GPU [^5] We used a single GPU to train the models with training sample sizes less than or equal to $100\,\rm K$, and multiple GPUs for models trained on more data. and a patience of 10 epochs. For training samples with sizes up to $100\,\rm K$ jets, the validation set was the same size as the training set. For larger samples, we used validation set sizes of $200\,\rm K$ jets for $1\,\rm M$ training samples and $500\,\rm K$ jets for $10\,\rm M$ training samples.

### v2

@@ Abstract
Foundation models are deep learning models pre-trained on large amounts of data which are capable of generalizing to multiple datasets and/or downstream tasks. This work demonstrates how data collected by the CMS experiment at the Large Hadron Collider can be useful in pre-training foundation models for HEP. Specifically, we introduce the AspenOpenJets dataset, consisting of approximately 178M high $p_{T}$ jets derived from CMS 2016 Open Data. We show how pre-training the OmniJet-$\alpha$ foundation model on AspenOpenJets improves performance on generative tasks with significant domain shift: generating boosted top and QCD jets from the simulated JetClass dataset. In addition to demonstrating the power of pre-training of a jet-based foundation model on actual proton-proton collision data, we provide the ML-ready derived AspenOpenJets dataset for further public use.
@@ I Introduction
While particle physics has long used machine learning techniques and is leading the way in adopting modern deep learning methods, the development and application of powerful foundation models – which have transformed natural language processing and computer vision – is still in its early stages. For the purposes of this work, a foundation model will be a model that has been pre-trained on a large amount of data for a specific task, and that can be used for different tasks downstream bommasani2022opportunities. The pre-training thus serves as a foundation, upon which the downstream tasks can be built. Well-known foundation models include large language models (LLMs) such as GPT-3 brown2020language, BERT devlin2019bert and LLaMA touvron2023llama, image models such as Imagen saharia2022photorealistic, and mixed-modality models such as CLIP radford2021clip and ALIGN DBLP:journals/corr/abs-2102-05918.
@@
There are several advantages to the use of foundation models. One example is the potential to boost the performance on small datasets. Having access to a model pre-trained on a larger, similar dataset provides a ‘head start’ for the downstream task, allowing the model to focus on the intricacies of the smaller dataset rather than having to rediscover the general structure of the data. In addition, a foundation model can save both human and computational resources. While pre-training may be a resource-intensive task, the downstream models would require less training, less data, and less time spent on optimization compared to what would be required if these models were built from scratch. Another interesting aspect is that a trained foundation model could enable implicit sharing of data and resources between different parts of the research community, a sharing that would not otherwise be possible.
@@
Foundation models are poised to play a central role in particle physics, especially in the context of experiments at the Large Hadron Collider (LHC). They are able to process the huge amounts of complex and diverse data generated by the LHC and have the potential to integrate multiple modalities - such as images, high-level event features or particle four-vectors - allowing more comprehensive data analyses. Current work on foundation models for particle physics has mainly focused on different event generation and classification tasks Kishimoto:2023cys; Qu:2022mxj; Heinrich:2024sbg; Birk:2024knn; Harris:2024sra; Mikuni:2024qsr, investigating pre-training strategies, encoding schemes and different architectures, and quantifying the transfer learning and generalization capabilities of these models.
@@
In this paper, we explore the potential of using large quantities of Open Data from the LHC for pre-training foundation models. The LHC experiments have made significant amounts of data publicly available CMS:2011A; CMS:2012A; CMS:2016G; CMS:2016H; ATLAS:OPEN. Foundation models can leverage the LHC open data for training and fine-tuning, potentially integrating different data modalities and data from different experiments. Furthermore, training foundation models on real data will result in a more accurate representation of particle physics processes than training on simulations. Finally, foundation models trained on a large variety of actual LHC data will help to promote transparency and collaboration in particle physics research.
Here we present the AspenOpenJets (AOJ) dataset and study the transfer learning of foundation models obtained with this dataset. The AOJ dataset has been constructed from the CMS 2016 ‘JetHT’ open data CMS:2016G; CMS:2016H and consists of $178\,\rm M$ jets, making it the largest ML-ready dataset for LHC applications. Jets are important objects at the LHC, consisting of collimated sprays of particles. Previous open data studies of CMS jets have explored the substructure and energy flow of the jets Larkoski:2017bvj; Tripathee:2017ybi; Komiske:2019jim and the usefulness of open data for quark/gluon tagging Dolan:2023abg, but this is the first study to prepare and make public a form of the CMS open data that allows easy use of the dataset for phenomenological studies, especially for training machine learning models.
@@
As an initial proof-of-concept study, we pre-train the OmniJet-$\alpha$  Birk:2024knn foundation model on the AOJ dataset, which is expected to mostly contain QCD jets, and then fine-tune it to generate jets of different types from the JetClass JetClass dataset of simulated jets. We do two such fine-tunings: first for JetClass QCD jets, which occupy a different kinematic region compared to the jets in AOJ, and then for top jets, which differ significantly in several features. We show that pre-training on AOJ indeed benefits the downstream task, making it easier for the model to generate these types of jets with fewer training examples compared to training from scratch.
@@
The utility of CMS open data for pre-training has been previously explored in Kishimoto:2023cys, using 1M single-lepton events at the object level which were collected by CMS in 2015. Here we enlarge the scope and complexity by several orders of magnitude to encompass 178M jet events at the full constituent level, that is, including not only the high-level features of the jets themselves, but also the features of all the particles inside the jet. Furthermore, the authors of Kishimoto:2023cys used event selections to shape the open data into a similar topology to the intended target simulated data. In our work, no such shaping is needed.
@@ II Data
Figure 1: Comparison of AspenOpenJets (AOJ) and QCD JetClass distributions for particle and jet level observables. Here $d_{0}$ and $d_{z}$ represent the transverse and longitudinal impact parameter respectively, with $\sigma_{d_{0}}$ and $\sigma_{d_{z}}$ being their uncertainties, and $\tau_{21}$ and $\tau_{32}$ are N-subjettiness variables. The last panel shows the average number of jet constituents for each particle species, accompanied by the corresponding $1\sigma$ standard deviation bands.
@@ II.1 CMS open data
The CMS experiment CMS:2008xjf is a large general-purpose detector at the LHC. The CMS collaboration has recently released 16.4 fb-1 of data from their 2016 runs (G and H) CMS:2016G; CMS:2016H and a corresponding set of simulations. In contrast to previous data releases CMS:2010B; CMS:2011A; CMS:2012A, this data has been provided in the simplified MINIAOD format Petrucciani:2015gjw, as well as the even further simplified NANOAOD format Peruzzi:2020. In these simplified analysis formats, the central CMS reconstruction steps have been applied, discarding much of the low-level content of the event in favor of a simplified description of the high-level objects (jets, electrons, muons, etc.) commonly used in data analysis. These formats are therefore significantly easier to use for those outside of the collaboration.
@@
CMS uses a particle-flow algorithm CMS:2017yfk, which combines the information from the different detector systems to reconstruct and identify each individual particle in an event. Particles are classified as either muons, electrons, photons, charged hadrons or neutral hadrons. Jets are clustered from the particle-flow candidates in an event using the anti-$k_{t}$ jet finding algorithm Cacciari:2008gp; Cacciari:2011ma. The typical large-radius jet size used in CMS for jet substructure searches and measurements is $R=0.8$, commonly referred to as AK8 jets. For AK8 jets the pileup-per-particle identification algorithm (PUPPI) Sirunyan:2020foa; Bertolini:2014bba is used to mitigate the effect of pileup. The PUPPI algorithm assigns a weight to each particle based on the probability that it originates from the primary vertex or from pileup. The momentum of AK8 jets is determined from the weighted vector sum of all particles inside the jet. Additional corrections are applied based on simulation and data-driven measurements to better calibrate the measured jet momentum to the true momentum of the particle from the hard process CMS:2016lmd.
II.2 Extracting Aspen Open Jets
The NANOAOD format does not natively store information of the jet constituents. However, an extended version (PFNANO), which does include this information, can be easily produced from MINIAOD using tools provided by CMS PFNANO. We therefore, for all data and simulations samples used, start with the MINIAOD format, process it to PFNANO and then apply our selections to select the jets of interest for our dataset.
In order to construct the largest sample of jets possible, we consider all events from the ‘JetHT’ CMS:2016G; CMS:2016H Run-2 dataset released by CMS to this date. The ‘JetHT’ dataset includes all events that have been selected by one of the HLT triggers related to jet momenta or total event hadronic activity (HT). Events are required to either have a high scalar sum of jet transverse momenta, or contain at least one or two jets with high energy. For the exact details of all different trigger paths used, see Refs. CMS:2016G; CMS:2016H. Because we do not specify a specific trigger, the sample consists of events from an inclusive ‘or’ of any of the available triggers of this dataset. We note that this choice works well for building a large dataset for the purposes of foundation model training, but may introduce subtle kinematic biases resulting from the combined thresholds of the different triggers, which would complicate its use in a physics-based analysis.
@@ II.2 Extracting Aspen Open Jets
Applying these selections to the CMS open data yields a dataset of 178 million jets, which we call the AspenOpenJets (AOJ). Due to the large cross section of QCD processes, the vast majority of these jets are expected to originate from light quarks and gluons. A small fraction of the jets are expected to originate from the decays of boosted heavy resonances. A rough estimate of the number of boosted fully merged jets produced by W, Z, top, and Higgs bosons in the $b\bar{b}$ decay mode are estimated with Monte Carlo (MC) simulations provided by CMS CMS:2016MC, with the remainder assumed to originate from QCD. The numbers are reported in Table 1. More details on the MC samples used are given in Appendix A.
@@
For each jet we store its transverse momentum ($p_{T}$), pseudorapidity ($\eta$), and azimuthal angular coordinate ($\phi$). We also store its mass, groomed with the softdrop algorithm softdrop as computed within the CMS reconstruction. Up to 150 constituents of the jet are stored. For each constituent, its 4-momentum is stored in the format $(p_{x},p_{y},p_{z},E)$. We additionally store its transverse impact parameter ($d_{0}$) and longitudinal impact parameter ($d_{z}$) with their uncertainties, the charge of the candidate, its particle-ID (PID) in the PDG format [^2] Note that neutral hadrons $({\rm h}^{0})$ are assigned the ${\rm PID}=130$ of the neutral kaon $K_{L}^{0}$ while positively/negatively charged hadrons $({\rm h}^{\pm})$ are assigned ${\rm PID}=\pm 211$ of the charged pion. ParticleDataGroup:2024cfk and its weight from the PUPPI algorithm. We also include additional jet substructure quantities computed within the CMS reconstruction, including the number of constituents in the jet, $N$-subjettiness variables nsubjettiness, various jet-tagging observables from the CMS implementation of ParticleNet particlenet; particlenet-cms and a regression of the jet mass from ParticleNet particlenet-cms.
@@
II.3 JetClass
The JetClass JetClass dataset is based on simulations and was first introduced in Qu:2022mxj. It contains both particle-level and jet-level information for $125\,\rm M$ simulated jets, distributed over 10 different jet types and divided into a training ($100\,\rm M$), validation ($5\,\rm M$) and test ($20\,\rm M$) set. The simulation is performed with MadGraph5_aMC@NLO Alwall:2014hca for the hard process, followed by Pythia Sjostrand:2014zea for parton showering and hadronization, and Delphes deFavereau:2013fsa with the CMS detector CMS:2008xjf card for the detector response. Jets are required to have a transverse momentum of $500\,\mathrm{GeV}<p_{T}<1000\,\mathrm{GeV}$ and a pseudorapidity $\eta<2$, which is more restrictive than the criteria used for AOJ. We overlay the JetClass QCD particle- and jet-level distributions (blue histograms) in fig. 1 to facilitate easy comparison with the corresponding AOJ distributions.
@@ III.1 Method
The foundation model chosen for this work is OmniJet-$\alpha$  Birk:2024knn, a GPT-style Radford2018ImprovingLU model that predicts the next token in a sequence. The features of the jet constituents [^3] The jet constituents are $p_{T}$ ordered and up to the first 128 constituents are considered when training the models. are first converted into integer tokens using a VQ-VAE oord2018neural; bao2022beit; Heinrich:2024sbg; huh2023straightening model, representing jets as sequences of tokens. The backbone of OmniJet-$\alpha$ is trained to predict the next token in these sequences. After pre-training, the model can generate new sequences autoregressively, which are then decoded by the frozen VQ-VAE back into physical space. Additionally, the pre-trained backbone weights can be loaded and used for other tasks, such as jet classification. In this work, we focus on dataset transfer learning for jet generation and leave the exploration of task switching for future studies.
@@
The features used for training are the constituent kinematics ($p_{T}$, $\Delta\eta$, $\Delta\phi$), where $\Delta\eta=\eta-\eta_{\text{jet}}$ and $\Delta\phi=\phi-\phi_{\text{jet}}$ represent the relative coordinates with respect to the jet axis. Before tokenization, these features were preprocessed using standardization [^4] Standardization is a common technique used to bring feature values into a manageable range for neural networks. By subtracting the mean and dividing by the standard deviation, the data is transformed to having a mean of zero and standard deviation of one.. The AOJ dataset was tokenized using a codebook size of 8192 tokens, employing VQ-VAE hyperparameters as outlined in Birk:2024knn, with two modifications: the number of attention heads was reduced from 8 to 4 without sacrificing performance, and the AdamW optimizer was replaced with the Ranger optimizer Ranger. A larger codebook with $\sim\!32\,\rm K$ tokens was also tested but produced results comparable to the $8\,\rm K$ configuration. Therefore, we only present results for the $8\,\rm K$ tokenization. The backbone model was trained on the full $178\,\rm M$ AOJ dataset with hyperparameters as specified in Birk:2024knn. The model was trained for 900K steps on 12 GPUs with a batch size of 256 per GPU, and the final state at the end of the training was selected as the backbone model.
@@
To evaluate the transfer learning capabilities of OmniJet-$\alpha$ , we devised a downstream task to test its ability to generate jets of a different type than those it was pre-trained on, and to adapt to datasets with different kinematic cuts. The fine-tuning process involved loading the pre-trained backbone weights and retraining it using jets from the JetClass JetClass dataset. To systematically evaluate the performance of the model, we trained it on six datasets with $10^{2}$, $10^{3}$, $10^{4}$, $10^{5}$, $10^{6}$, $10^{7}$ jets to study the effect of dataset size $D$, and fine-tuned the model separately on QCD jets ($q/g$) and top-quark jets ($t\to bqq^{\prime}$) to evaluate its adaptability to different jet substructure types. For each configuration, the model with the lowest validation loss during fine-tuning was selected for jet generation. To provide a baseline for comparison, we used a separate VQ-VAE tokenizer and trained generative models with the same backbone architecture from scratch using randomly initialized weights. This VQ-VAE is the one from Ref. Birk:2024knn that was trained on all 10 classes of the JetClass dataset. The generative models were trained on the tokenized QCD and top-quark samples. This approach allowed us to quantify the performance improvement achieved by pre-training on the AOJ dataset.
@@
Both training strategies used a batch size of 256555For training samples of 100 jets, the batch size was reduced to 100. per GPU [^6] We used a single GPU to train the models with training sample sizes less than or equal to $100\,\rm K$, and multiple GPUs for models trained on more data. and a patience of 10 epochs. For training samples with sizes up to $100\,\rm K$ jets, the validation set was the same size as the training set. For larger samples, we used validation set sizes of $200\,\rm K$ jets for $1\,\rm M$ training samples and $500\,\rm K$ jets for $10\,\rm M$ training samples.